# Quasilinear problems with mixed local-nonlocal operator and concave-critical nonlinearities: Multiplicity of positive solutions

### v1

@@ Abstract.
where $\Omega\subset\mathbb{R}^{N}$ is a bounded open set, $\varepsilon\in(0,1]$, $0<s<1<q<p<N$, and $p^{*}=\frac{Np}{N-p}$, and $\lambda\in\mathbb{R}$ is a parameter. For $\lambda\leq 0$, we show that ($\mathcal{P}_{\lambda,\varepsilon}$) has no nontrivial solution. For $\lambda>0$, we prove Ambrosetti-Brezis-Cerami type results. In particular, we prove the existence of $\Lambda_{\varepsilon}$ such that ($\mathcal{P}_{\lambda,\varepsilon}$) has a positive minimal solution for $0<\lambda<\Lambda_{\varepsilon}$, a positive solution for $\lambda=\Lambda_{\varepsilon}$ and no positive solution for $\lambda>\Lambda_{\varepsilon}$. We also prove the existence of $0<\lambda^{\#}\leq\Lambda_{\varepsilon}$ such that ($\mathcal{P}_{\lambda,\varepsilon}$) has at least two positive solutions for $\lambda\in(0,\lambda^{\#})$ provided $\varepsilon$ small enough. This extends the recent result of Biagi and Vecchi (Nonlinear Anal. 256 (2025), 113795), Amundsen, et al. (Commun. Pure Appl. Anal., 22(10):3139–3164, 2023) from $p=2$ to the general $1<p<N$. Additionally, it extends the classical result of Azorero and Peral (Indiana Univ. Math. J., 43(3):941–957, 1994) to the mixed local-nonlocal quasilinear problems. Moreover, our results complements the multiplicity results for nonnegative solutions in da Silva, et al. (J. Differential Equations, 408:494–536, 2024).
@@ 1. Introduction
  • a second solution for any $\lambda\in(0,\Lambda)$.
@@ 3. Existence and nonexistence of solutions
Let $q\in(1,p)$ and $\varepsilon\in(0,1]$. Then there exist $r_{0},\,\lambda^{\#}>0$ independent of $\varepsilon$, such that for every $0<\lambda<\lambda^{\#}$, ($\mathcal{P}_{\lambda,\varepsilon}$) has a positive solution $u_{\lambda,\varepsilon}$ in the interior of $B_{r_{0}}$.
@@
By the Sobolev embedding $X_{0}\hookrightarrow L^{p^{*}}(\Omega)$, we can choose $r_{0}>0$ and $\delta_{0}>0$ small enough such that
$\frac{1}{p}\rho_{\varepsilon}(u)^{p}-\frac{1}{p^{\ast}}\|u_{+}\|_{p^{\ast}}^{p^{\ast}}\geq\frac{1}{p}\rho_{\varepsilon}(u)^{p}-C_{1}\rho_{\varepsilon}(u)^{p^{\ast}}\geq\left\{\begin{array}[]{ll}0,&\forall\,u\in B_{r_{0}};\\ 2\delta_{0},&\forall\,u\in\partial B_{r_{0}}.\\ \end{array}\right.$ (3.4)
Let $I_{\lambda,\varepsilon}$ be as defined in (2.4). Therefore, $I_{\lambda,\varepsilon}(u)\geq-\frac{\lambda}{q}\|u_{+}\|_{q}^{q}\geq-\frac{\lambda C}{q}r_{0}^{q}>-\infty$ for every $u\in B_{r_{0}}$. Thus $I_{\lambda,\varepsilon}$ is bounded from below on $B_{r_{0}}$. Define, $c_{\lambda,\varepsilon}:=\inf_{B_{r_{0}}}I_{\lambda,\varepsilon}(u).$ For any $v\neq 0$ in $X_{0}$ and $t>0$,
@@
It is easy to see that for small $t$, $I_{\lambda,\varepsilon}(tv)<0$. Hence $-\infty<c_{\lambda,\varepsilon}<0$. Consider a minimizing sequence $u_{n,\varepsilon}\in B_{r_{0}}$ i.e. $I_{\lambda,\varepsilon}(u_{n,\varepsilon})\to c_{\lambda,\varepsilon}\text{ as}n\to\infty.$ We claim that $\rho_{\varepsilon}(u_{n,\varepsilon})\leq{r_{0}}-\varepsilon_{0}$ for every $n$ and for some $\varepsilon_{0}$ independent of $n$. If not, $\rho_{\varepsilon}(u_{n,\varepsilon})\to{r_{0}}$. Thus, by Sobolev inequality,
$c_{\lambda,\varepsilon}=\lim_{n\to\infty}I_{\lambda,\varepsilon}(u_{n,\varepsilon})\geq\lim_{n\to\infty}\left(\frac{1}{p}\rho_{\varepsilon}(u_{n,\varepsilon})^{p}-C_{1}\rho_{\varepsilon}(u_{n,\varepsilon})^{p^{\ast}}-\lambda C_{2}\rho_{\varepsilon}(u_{n,\varepsilon})^{q}\right)\geq 2\delta_{0}-\lambda C_{2}r_{0}^{q}.$
Choosing $\lambda_{\ast\ast}>0$ such that for every $0<\lambda<\lambda_{\ast\ast}$ it holds $2\delta_{0}-\lambda C_{2}r_{0}^{q}>\delta_{0}$. Thus we get
@@
which yields a contradiction and hence $u_{n,\varepsilon}\in B_{r_{0}-\varepsilon_{0}}$ for every $n$. Let $0<\varepsilon_{1}<\varepsilon_{0}$, then $B_{r_{0}-\varepsilon_{0}}\subset B_{r_{0}-\varepsilon_{1}}$. Applying Ekeland’s variational principle on the complete metric space $B_{r_{0}-\varepsilon_{1}}$ (w.r.t Euclidean metric), we get
@@
$I_{\lambda,\varepsilon}(u_{n,\varepsilon})\leq I_{\lambda,\varepsilon}(v)+\frac{1}{n}\rho_{\varepsilon}(u_{n,\varepsilon}-v),\;\forall\,v\in B_{r_{0}-\varepsilon_{1}},\,v\neq u_{n,\varepsilon}.$
@@
Let $w\in X_{0}$ and $t>0$ be such that $\rho_{\varepsilon}(w)=1$ and $v=u_{n,\varepsilon}+tw\in B_{r_{0}-\varepsilon_{1}}$. Therefore, using the above relation we get
@@
Hence, for any $0<\lambda<\lambda^{\#}:=\min\{\lambda^{\ast},\lambda_{\ast\ast}\},$ there exists $u_{\lambda,\varepsilon}\in B_{r_{0}-\varepsilon_{0}}$ such that $u_{n,\varepsilon}\to u_{\lambda,\varepsilon}$ in $X_{0}$, $I_{\lambda,\varepsilon}(u_{\lambda,\varepsilon})=c_{\lambda,\varepsilon}$ and $u_{\lambda,\varepsilon}$ is a critical point of $I_{\lambda,\varepsilon}$. Hence $u_{\lambda,\varepsilon}$ is a positive solution of ($\mathcal{P}_{\lambda,\varepsilon}$). ∎
@@ 4. Multiplicity of positive solutions
    If there exists $r_{0}\in(0,1)$, independent of $\varepsilon$, such that $\rho_{\varepsilon}(u)\leq r_{0}$, then
@@
(b) To show (b), we return to (4.10). Using the Sobolev inequality and the fact that $\|\nabla u\|_{p}\leq\rho_{\varepsilon}(u)\leq r_{0}$,
$\|\varphi(u)\|_{p^{\ast}}^{p}\leq\frac{2(\lambda_{0}+1)}{S_{0}^{\frac{p^{\ast}+p}{p}}}\beta_{1}^{p-1}(R^{p^{\ast}-1}+1)\,r_{0}^{p^{\ast}}.$
Taking $T\to\infty$, we see that $A_{1}\leq C$ is independent of $\varepsilon$. Hence, $\lVert u\rVert_{\infty}\leq C$ where $C$ is independent of $\varepsilon$.
@@
Since in Proposition 4.1, we have chosen $\lambda_{0}$ arbitrarily and then fixed it, in view of Proposition 3.4, we can choose $\lambda_{0}=\lambda^{\#}$ so that for every $0<\lambda<\lambda^{\#}$ we have $\|u_{\lambda,\varepsilon}\|_{\infty}\leq C$ where $C$ is independent of $\varepsilon$.

### v2

@@ Abstract.
where $\Omega\subset\mathbb{R}^{N}$ is a bounded open set, $\varepsilon\in(0,1]$, $0<s<1<q<p<N$, and $p^{*}=\frac{Np}{N-p}$, and $\lambda\in\mathbb{R}$ is a parameter. For $\lambda\leq 0$, we show that ($\mathcal{P}_{\lambda,\varepsilon}$) has no nontrivial solution. For $\lambda>0$, we prove Ambrosetti-Brezis-Cerami type results. In particular, we prove the existence of $\Lambda_{\varepsilon}$ such that ($\mathcal{P}_{\lambda,\varepsilon}$) has a positive minimal solution for $0<\lambda<\Lambda_{\varepsilon}$, a positive solution for $\lambda=\Lambda_{\varepsilon}$ and no positive solution for $\lambda>\Lambda_{\varepsilon}$. We also prove the existence of $0<\lambda^{\#}\leq\Lambda_{\varepsilon}$ such that ($\mathcal{P}_{\lambda,\varepsilon}$) has at least two positive solutions for $\lambda\in(0,\lambda^{\#})$ provided $\varepsilon$ small enough. This extends the recent result of Biagi and Vecchi (Nonlinear Anal. 256 (2025), 113795), Amundsen, et al. (Commun. Pure Appl. Anal., 22(10):3139–3164, 2023) from $p=2$ to the general $1<p<N$. Additionally, it extends the classical result of Azorero and Peral (Indiana Univ. Math. J., 43(3):941–957, 1994) to the mixed local-nonlocal quasilinear problems. Moreover, our results complement the multiplicity results for nonnegative solutions in da Silva, et al. (J. Differential Equations, 408:494–536, 2024).
@@ 1. Introduction
  • a second solution for any $\lambda\in(0,\Lambda)$, provided $r\leq 2^{*}$.
@@ 3. Existence and nonexistence of solutions
Let $q\in(1,p)$ and $\varepsilon\in(0,1]$. Then there exist $r_{0}>0$ independent of $\varepsilon$ such that for every $0<r\leq r_{0}$, there exists $\lambda^{\#}>0$, independent of $\varepsilon$, such that for every $0<\lambda<\lambda^{\#}$, ($\mathcal{P}_{\lambda,\varepsilon}$) has a positive solution $u_{\lambda,\varepsilon}$ in the interior of $B_{r}$.
@@
By the Sobolev embedding $X_{0}\hookrightarrow L^{p^{*}}(\Omega)$, we can choose $r_{0}>0$ and $\delta_{0}>0$ small enough such that, for every $0<r\leq r_{0}$,
$\frac{1}{p}\rho_{\varepsilon}(u)^{p}-\frac{1}{p^{\ast}}\|u_{+}\|_{p^{\ast}}^{p^{\ast}}\geq\frac{1}{p}\rho_{\varepsilon}(u)^{p}-C_{1}\rho_{\varepsilon}(u)^{p^{\ast}}\geq\left\{\begin{array}[]{ll}0,&\forall\,u\in B_{r};\\ 2\delta_{0},&\forall\,u\in\partial B_{r}.\\ \end{array}\right.$ (3.4)
Let $I_{\lambda,\varepsilon}$ be as defined in (2.4). Therefore, $I_{\lambda,\varepsilon}(u)\geq-\frac{\lambda}{q}\|u_{+}\|_{q}^{q}\geq-\frac{\lambda C}{q}r_{0}^{q}>-\infty$ for every $u\in B_{r}$. Thus $I_{\lambda,\varepsilon}$ is bounded from below on $B_{r}$. Define, $c_{\lambda,\varepsilon}:=\inf_{B_{r_{0}}}I_{\lambda,\varepsilon}(u).$ For any $v\neq 0$ in $X_{0}$ and $t>0$,
@@
It is easy to see that for small $t$, $I_{\lambda,\varepsilon}(tv)<0$. Hence $-\infty<c_{\lambda,\varepsilon}<0$. Consider a minimizing sequence $u_{n,\varepsilon}\in B_{r}$ i.e. $I_{\lambda,\varepsilon}(u_{n,\varepsilon})\to c_{\lambda,\varepsilon}\text{ as}n\to\infty.$ We claim that $\rho_{\varepsilon}(u_{n,\varepsilon})\leq{r}-\varepsilon_{0}$ for every $n$ and for some $\varepsilon_{0}$ independent of $n$. If not, $\rho_{\varepsilon}(u_{n,\varepsilon})\to{r}$. Thus, by Sobolev inequality,
$c_{\lambda,\varepsilon}=\lim_{n\to\infty}I_{\lambda,\varepsilon}(u_{n,\varepsilon})\geq\lim_{n\to\infty}\left(\frac{1}{p}\rho_{\varepsilon}(u_{n,\varepsilon})^{p}-C_{1}\rho_{\varepsilon}(u_{n,\varepsilon})^{p^{\ast}}-\lambda C_{2}\rho_{\varepsilon}(u_{n,\varepsilon})^{q}\right)\geq 2\delta_{0}-\lambda C_{2}r^{q}.$
Choosing $\lambda_{\ast\ast}>0$ such that for every $0<\lambda<\lambda_{\ast\ast}$ it holds $2\delta_{0}-\lambda C_{2}r^{q}>\delta_{0}$. Thus we get
@@
which yields a contradiction and hence $u_{n,\varepsilon}\in B_{r-\varepsilon_{0}}$ for every $n$. Let $0<\varepsilon_{1}<\varepsilon_{0}$, then $B_{r-\varepsilon_{0}}\subset B_{r-\varepsilon_{1}}$. Applying Ekeland’s variational principle on the complete metric space $B_{r-\varepsilon_{1}}$ (w.r.t Euclidean metric), we get
@@
$I_{\lambda,\varepsilon}(u_{n,\varepsilon})\leq I_{\lambda,\varepsilon}(v)+\frac{1}{n}\rho_{\varepsilon}(u_{n,\varepsilon}-v),\;\forall\,v\in B_{r-\varepsilon_{1}},\,v\neq u_{n,\varepsilon}.$
@@
Let $w\in X_{0}$ and $t>0$ be such that $\rho_{\varepsilon}(w)=1$ and $v=u_{n,\varepsilon}+tw\in B_{r-\varepsilon_{1}}$. Therefore, using the above relation we get
@@
Hence, for any $0<\lambda<\lambda^{\#}:=\min\{\lambda^{\ast},\lambda_{\ast\ast}\},$ there exists $u_{\lambda,\varepsilon}\in B_{r-\varepsilon_{0}}$ such that $u_{n,\varepsilon}\to u_{\lambda,\varepsilon}$ in $X_{0}$, $I_{\lambda,\varepsilon}(u_{\lambda,\varepsilon})=c_{\lambda,\varepsilon}$ and $u_{\lambda,\varepsilon}$ is a critical point of $I_{\lambda,\varepsilon}$. Hence $u_{\lambda,\varepsilon}$ is a positive solution of ($\mathcal{P}_{\lambda,\varepsilon}$). ∎
@@ 4. Multiplicity of positive solutions
    There exists $r_{0}\in(0,1)$, independent of $\varepsilon$, such that if $\rho_{\varepsilon}(u)\leq r_{0}$, then
@@
(b) To show (b), we return to (4.8): choosing $R=1$, using the Sobolev inequality and the fact that $\|\nabla u\|_{p}\leq\rho_{\varepsilon}(u)\leq r_{0}$, we get
$\int_{\Omega}(\varphi(u))^{p}|u|^{p^{*}-p}\,{\rm d}x$ $\leq\int_{\Omega}|u|^{p^{\ast}}\,{\rm d}x+\left(\int_{\Omega}(\varphi(u))^{p^{\ast}}\,{\rm d}x\right)^{\frac{p}{p^{\ast}}}\left(\int_{\{|u|\geq 1\}}|u|^{p^{\ast}}\,{\rm d}x\right)^{\frac{p^{\ast}-p}{p^{\ast}}}$
$\leq\int_{\Omega}|u|^{p^{\ast}}\,{\rm d}x+\left(\int_{\Omega}(\varphi(u))^{p^{\ast}}\,{\rm d}x\right)^{\frac{p}{p^{\ast}}}\left(\int_{\Omega}|u|^{p^{\ast}}\,{\rm d}x\right)^{\frac{p^{\ast}-p}{p^{\ast}}}$
$\leq\int_{\Omega}|u|^{p^{\ast}}\,{\rm d}x+\left(\int_{\Omega}(\varphi(u))^{p^{\ast}}\,{\rm d}x\right)^{\frac{p}{p^{\ast}}}S_{0}^{\frac{p-p^{*}}{p}}r_{0}^{p^{*}-p}$
Choosing $r_{0}>0$ such that
$S_{0}^{\frac{p-p^{*}}{p}}r_{0}^{p^{*}-p}\leq\frac{1}{2C_{0}\beta_{1}^{p-1}}.$
Proceeding as before and taking $T\to\infty$, we see that $A_{1}\leq C$ is independent of $\varepsilon$. Hence, $\lVert u\rVert_{\infty}\leq C$ where $C$ is independent of $\varepsilon$.
@@
Since in Proposition 4.1, we have chosen $\lambda_{0}$ arbitrarily and then fixed it, in view of Proposition 3.4, we can choose $\lambda_{0}=\lambda^{\#}$ and $r_{0}>0$ so that for every $0<\lambda<\lambda^{\#}$ we have $u_{\lambda,\varepsilon}$ lies in the interior of $B_{r_{0}}$ and $\|u_{\lambda,\varepsilon}\|_{\infty}\leq C$ where $C$ is independent of $\varepsilon$.